OpenStudy (anonymous):

What is the value of the expression (–2)^5? A. –32 B. –10 C. 32 D. 10

OpenStudy (anonymous):

-32 because it's an odd power.

OpenStudy (unklerhaukus):

\[\large a^n=\overbrace{a\times a\times a\cdots\times a}^{n\text{ times}}\] \[(-2)^5=\overbrace{-2\times-2\times-2\times-2\times-2}^{\text{5 times}}\]

OpenStudy (anonymous):

thanks both of you